OpenStudy (anonymous):

A 0.12 kg ball of dough is thrown straight up into the air with an initial speed of 11 m/s. The acceleration of gravity is 9.8 m/s2 What is its momentum halfway to its maxi- mum height? Answer in units of kgm/s

OpenStudy (anonymous):

Not certain if correct, but hope it provides something.

OpenStudy (anonymous):

First, find the half of the maximum that the ball will reach. Use the formula \[\frac{ vf -vi }{ a }=t\] to find the time.\[vf=0m/s, vi=11m/s, a=-9.8m/s ^{2}\] \[\frac{ 0-11 }{-9.8 }=\frac{ 11 }{ 9.8}s\] Now use this formula \[vt+\frac{ 1 }{ 2 }a ^{2}=s\] then get \[11\times \frac{ 11 }{ 9.8 }-\frac{ 1 }{ 2 }\times9.8\times \left( \frac{ 11 }{ 9.8 } \right)^{2}=\frac{ 1 }{ 2 }\times \frac{ 11^{2} }{ 9.8}\] And \[\frac{ 1 }{ 2 }s =\frac{ 1 }{ 4 }\times \frac{ 11^{2} }{ 9.8 }\] Second, find the time that the ball spend to reach the half of the the maximum it will reach.\[vf ^{2}-vi ^{2}=2gs\] \[vf ^{2}=2\times \left( -9.8 \right)\times \frac{ 1 }{ 4 }\times \frac{ 11^{2} }{ 9.8 }+11^{2}\] \[vf ^{2}=\frac{ 1 }{ 2 }\times11^{2}\] \[vf =11\sqrt{\frac{ 1 }{ 2 }}m/s\] As p(momentum)=mv, the momentum of the ball is \[0.12\times11\sqrt{\frac{ 1 }{ 2 }}=1.32\sqrt{\frac{ 1 }{ 2 }}kg \times m/s\]

OpenStudy (anonymous):

You should be able to skip solving for any sort of time just by using the 4th equation of motion. Use V = final velocity = 0 m/s Vi = 11 m/s h = max height \[ v^2 = v_i^2 + 2ah \\ -v_i^2 = -2gh \\ h = \frac{v_i^2}{2g} \] Then just use the same equation with a different final velocity and new height; Vi is the same, V is the unknown, and h/2 is half the max height by definition. \[ v_{(h/2)}^2 = v_i^2 + 2a{\frac{h}{2}} \\ v_{(h/2)}^2 = v_i^2 - g \left( \frac{v_i^2}{2g} \right) \\ v_{(h/2)}^2 = \frac{v_i^2}{2} \\ v_{(h/2)} = v_i \frac{\sqrt2}{2} \] then momentum \[ p_{(h/2)}=mv_{(h/2)} \\ p_{(h/2)} = mv_i \frac{\sqrt2}{2} \] Gets the same answer without the pesky time :)